if vector a= cos omega ti^+sin omega tj^ and b= cos omega t /2 i^+sin omega t /2 j^ are function of time then the value of t when they are orthogonal to

  1. if vector a= cos omega ti^+sin omega tj^ and b= cos omega t /2 i^+sin omega t /2 j^ are function of time then the value of t when they are orthogonal to

Grade:11

1 Answers

Khimraj
5 years ago
when they are orthogonal their dot product will be zero
a.b = coswt*coswt/2 + sinwt*sinwt/2 = cos(wt – wt/2) = coswt/2 = 0
so wt/2 = \pi /2
so t = \pi /\omega
Hope it clears........................